# Tightrope-walking rationality in action: Feyerabendian insights for the foundations of quantum mechanics

### v2

@@ 3 Realism vs instrumentalism, in the interpretation of QM
PKF’s target is the orthodoxy of the time, and its instrumentalist and positivist core. We note here, though, that we could direct a similar criticism, and with the same philosophical arguments, toward any realist interpretation of QM in which the physical reality we decide to preserve, despite the QM challenges, corresponds to an ontology built on the same premises of our observational language or our common sense, that is, upon reflection, on classical Newtonian mechanics. An example would be the same deBroglie-Bohm theory that was the only contender of the orthodox Copenhagen view at the time of PKF’s writing. The criticism would then extend to more modern perspectives, bohmian or not, aiming at re-establishing at the heart of fundamental physics a primitive ontology of particles moving in space and in presence of an absolute time, an ontology which is essentially Newtonian [22]. In connection to this, it is interesting to look at modern ”hidden variables” completions of QM, formalized in the mathematical framework of ”ontological models” [23], where quantum states are epistemic states (states of knowledge) over a hidden (from QM) set of ontic states of the physical system under consideration. These models (which at present can cover only ”fragments” of QM) are indeed forced by the lessons of QM (including both theoretical developments of the last decades and its observational successes) to look for a precise characterization of the ontic structure of the world (what is real) in terms much farther away from classical realism and a Newtonian picture than the early advocates of a realist alternative to the Copenhagen instrumentalism (and the Bohmians) would have liked (see, e.g. [24]).
@@
It is instructive, then, to read further PKF’s critical position about this dogmatic turn of empiricism, leading to the instrumentalist Copenhagen interpretation of QM, appreciating the anti-dogmatic spirit, but also putting the realist Bohmian or hidden variable position under the same critical lens (we could try to substitute any reference to Bohr or Copenhagen in this quote with a reference to Einstein or Bohm or other ”realist” proposal), to realize that the dogmatism attacked by PKF has spread beyond the Copenhagen interpretation:
Duality is regarded by Bohr and by his followers as an experimental fact which must not be tampered with and upon which all future reasoning about microphysical events is to be based. […] It follows, and this is Bohr’s contention, that it will hold for any theory into which Planck’s h enters in an essential way. Hence, any future microscopic theory will have to be descriptive of appearances only, it will contain irreducible probabilites, and it will have to work with commutation relations between variables which are only partly well defined and meaningful. The development of microphysics can only lead to greater indeterminancy. It will never again return to a state of affairs where we are able to give a complete, objective and deterministic description of the nature of physical systems and physical events. In the interest of economy of thought and effort, theories of this kind should therefore be forever excluded from consideration. […] The argument proceeds from what seems to be a mere truism; it proceeds from the assertion that, duality being an experimental fact, it must not be tampered with, but must be regarded as an unalterable basis for any further theoretical consideration. […] It is here, at the very beginning, that the position of the orthodox must be criticized. For what is regarded by them as a truism is neither correct nor reasonable; and their account of history, too, is at variance with the actual development. Things were just the other way round.
@@
The Galilean tradition, as we may call it, therefore proceeds from the very reasonable point of view that our ideas as well as our experiences (complicated experimental results included) may be erroneous, and that the latter give us at most an approximate account of what is going on in reality. Hence, within this tradition the condition to be satisfied by a future theory of the microcosm is not that it be simply compatible with duality and the other laws used in the above argument, but that it be compatible with duality to a certain degree of approximation which will have to depend on the precision of the experiments used for establishing the ’fact’ of duality. […] A completely analogous remark holds for the assertion that Planck’s constant will have to enter every microscopic theory in an essential way. After all, it is quite possible that this constant has meaning only under certain well-defined conditions (just as the density of a fluid, or its viscosity, or its diffusion constant can be meaningfully defined only for not too small a volume), and that the experiments we have made so far explore only part of these conditions. Obviously the invariance of $\hbar$ in all these experiments cannot be used as an argument against such a possibility. But if neither the constancy of h nor duality can be guaranteed to hold in new domains of research, then the whole argument is bound to break down: it does not guarantee the persistence of the familiar features of complementarity, of probabilistic laws, of quantum-jumps, of the commutation relations in future investigations. - [1], page 316-325
@@
In fact, conceptual revision is possible even about experimental facts which are taken to be exact and perfectly valid, and the historic example is again taken from Galileo’s story:
@@
Let us now assume, for the sake of argument, that a radically empiricistic point of view has been adopted, i.e. let us regard duality and the constancy of A as holding with absolute precision. Is the argument now valid? This at once introduces the second ’metaphysical’ assumption that is used by Bohr and his followers. According to this second assumption the classical concepts are the only concepts we have. As we cannot construct a theory or a description of fact out of concepts which we do not have, and as the classical concepts cannot any more be applied in an unrestricted way, we are stuck with the complementary mode of description. […] it is sufficient to point out that introducing a set of concepts is not something that occurs independently of and prior to the construction of theories. Concepts are introduced as part of a theoretical framework, they are not introduced by themselves. However, with respect to theories it must be asserted that man is not only capable of using them and the concepts which they embody for the construction of descriptions, experimental and otherwise, but that he is also capable of inventing them. How else could it have been possible, to mention only one example, to replace the Aristotelian physics and the Aristotelian cosmology by the new physics of Galileo and Newton? The only conceptual apparatus then available was the Aristotelian theory of change with its opposition of actual and potential properties, form and matter, the four causes and the like. This conceptual apparatus was much more general and universal than the physical theories of today for it contained a general theory of change, spatiotemporal and otherwise. It also seems to be closer to everyday thinking and was therefore more firmly entrenched than any succeeding physical theory, classical physics included. Within this tremendously involved conceptual scheme Galileo’s (or rather Descartes’) law of inertia does not make sense. Should, then, Galileo have tried to get on with the Aristotelian concepts as well as possible because these concepts were the only ones in actual use and as ’there is no use discussing what could be done if we were other [i.e. more ingenious] beings that we are?’ By no means! What was needed was not improvement, or delimitation of the Aristotelian concepts in order to ’make room for new physical laws’; what was needed was an entirely new theory. Now at the time of Galileo human beings were apparently able to do this extraordinary thing and become ’beings different from what they were before’ (and one should again realize that the conceptual change that was implied was much more radical than the conceptual change necessitated by the appearance of the quantum of action). Are there (apart from pessimism with respect to the abilities of contemporary physicists) any reasons to assume that what was possible in the sixteenth and seventeenth centuries will be impossible in the twentieth century? […] To sum up, although in reporting our experiences we make use, and must make use, of certain theoretical terms, it does not follow that different terms will not do the job equally well, or perhaps even better, because more coherently. And as our argument was quite general, it seems to apply to the classical concepts as well. This is where Bohr’s second argument comes in. […] The argument itself runs as follows: if classical physics is a universal theory (premise c) and has been used long enough (premise d), then all our experiences will be classical (premise b) and we shall therefore be unable to conceive any concepts which fall outside the classical scheme (premise a). Hence the invention of a new conceptual scheme which might enable us to circumvent duality is impossible. That there must be something amiss with the argument is seen from the fact that all the premises, except perhaps the first one, apply also to the Aristotelian theory of motion. […] However, the Aristotelian theory has been superseded by a very different conceptual apparatus. Clearly, this new conceptual apparatus was then not suggested by experience as interpreted in the Aristotelian manner and it was therefore a ’free mental creation’”. This refutes (a). That (b) needs modifying becomes clear when we consider that a scientist should always keep an open mind and that he should therefore always consider possible alternatives along with the theory that he is favouring at a certain moment. […] We have to conclude, then, that Bohr’s arguments against the possibility of alternatives to the point of view of complementarity are all inconclusive. - [1], page 316-325
@@
This result is exactly as it should be. Any restrictive demand with respect to the form and the properties of future theories can be justified only if an assertion is made to the effect that certain parts of the knowledge we possess are absolute and irrevocable. Dogmatism, however, should be alien to the spirit of scientific research, and this quite irrespective of whether it is now grounded upon ’experience’ or upon a different and more ’aprioristic’ kind of argument. - [1], page 316-325
@@
Having established that the distinction between theories and (empirical) facts, which is required to ground a philosophical distinction between realism and instrumentalism, is not an absolute one and that also the observational language is a theoretical one, PKF argues that the choice between realism and instrumentalism amounts to a strategic (and heuristic) choice in our pursue of a scientific understanding of nature. It has to do with which concepts (whether those grounding the (more) theoretical language or those grounding the (more) observational one) we decide to rely on when trying to improve our scientific understanding of the world. As with any strategic decision, we should always remain flexible and constantly adjust our strategies as we move forward and question the assumptions at the basis of them. We should, in other words, be ‘rationally tightrope-walking’.
@@
In other words, preferring realism over instrumentalism, or viceversa, is not something that can be decided on philosophical grounds (especially philosophical arguments regarding theory comparison, which are hindered by incommensurability), but by its results in facilitating progress, on a case by case basis and on a temporary basis, eve though the concrete implementation of a new general perspective can be difficult. And the fact that one cannot decide a priori in favor of one option over the other does not make the choice an arbitrary one either, because the consequences for scientific progress of taking one option or the other can be judged. In PKF’s words:
@@
the issue between positivism and realism is not a factual issue which can be decided by pointing to certain actually existing things, procedures, forms of language, etc., it is an issue between different ideals of knowledge. There seem to exist two objections against this characterization of the situation. The first objection is that it makes the resolution of the issue an arbitrary one. The second objection is that different ideals of knowledge cannot be realized equally easily. Taking the second objection first, we are prepared to admit that there may be psychological difficulties in inventing theories of a certain type, especially if metaphysical views are held which seem to recommend, radically different theories. However, a stronger variant of the second objection is frequently used. According to this stronger variant, all our theoretical knowledge is (uniquely) determined by the facts and cannot be chosen at will. Against this objection we repeat that what is determined by the ’facts’ is the acceptance (or rejection) of sentences which are already interpreted and which have been interpreted independently of the phenomenological character of what is observed. The impression that every fact suggests one and only one interpretation and that therefore our views are ’determined’ by the facts, this impression will arise only when (with respect to the language used) the relation of phenomenological adequacy is a one-one relation. As has been pointed out before, such a situation arises whenever a fairly general point of view was held long enough to influence our expectations, our language and thereby our perceptions, and when during that period no alternative picture was ever seriously considered. We may prolong such a situation either by explaining away adverse facts with the help of ad hoc hypotheses which are framed in terms of the points of view to be conserved; or by reducing more successful alternatives to ’instruments of prediction’ which, being devoid of descriptive meaning, cannot clash (in the phenomenological sense) with any experience; or by devising a criterion of significance according to which such alternatives are meaningless. […] But we can also choose the opposite procedure, i.e. we can take refutations seriously and regard alternative theories, in spite of their unusual character, as descriptive of really existing things, properties, relations etc. In short, although the truth of a theory may not depend upon us, its form (and the form of our theoretical knowledge in general) can always be arranged so as to satisfy certain demands. With this the second objection collapses. - [1], page 34-36
@@
Once more, the only real enemy is dogmatism, in all its forms and critical (self-)assessment is mandatory and should be a constant element of our investigations, whether we initially decide to proceed within a realist perspective or from an instrumentalist standpoint. One has to exercise tightrope-walking rationality, avoiding both a-critical realism and hypocritical instrumentalism, and keep going.
@@
Having accepted this fundamental anti-dogmatic stance, does PKF have arguments, however tentative and purely strategic in nature, for one or the other option? He does. It can be argued, he maintains, that realism tends to be less dogmatic than instrumentalism, since it does not take even empirical facts and observations for granted. Indeed, it treats all elements of scientific theories as subject to critical analysis, and leaves more room for metaphysical speculation. In this sense, it is always to be preferred [^6] It is interesting to notice that, starting from the same anti-dogmatic stance, and in fact based on a similar emphasis on the existence of alternative explanations for the same phenomena, one can also be led to prefer an instrumentalist rather than realist attitude. This is the case, for example, of K. Stanford’s version of instrumentalism [25].. This is a strong conclusion.
@@
Here are PKF’s words [^7] Incidentally, the following quote is also, at once, a justification of creativity in theory construction unhinged by empirical constraints and of non-empirical theory assessment, and also a reminder that empirical evidence remains the ultimate judge of of our theories.:
@@
But does the fact that the form of our knowledge can be adjusted to meet our demands, make the rejection of positivism an arbitrary act (first objection)? It does not, as we judge an ideal by the consequences which its realization may or may not imply. In what follows I want to discuss the consequences of the positivistic procedure and show why I think that realism is to be preferred. […] As opposed to positivism, a realistic position does not admit any dogmatic and incorrigible statement into the field of knowledge. Hence, also, our knowledge of what is observed is not regarded as unalterable and this in spite of the fact that it may have a counterpart in the phenomena them- selves. This means that at times interpretations will have to be considered which do not ’fit’ the phenomena and which clash with what is immediately given. Interpretations of this kind could not possibly emerge from close attention to the ’facts’. It follows that we need a non-observational source for interpretations. Such a source is provided by (metaphysical) speculation which is thus shown to play an important role within realism. However, the results of such speculation must be made testable, and having been transformed in this way they must be interpreted as descriptive of general features of the world (otherwise we are thrown back upon the old account of what is observed). This procedure (a) allows us to draw a clear boundary line between objective states of affair and the states of the observer, though it admits that we may be mistaken with regard to the exact position of the boundary line; it (b) is empirical in the sense that no dogmatic statements are allowed to become elements of knowledge; it (c) is liable to encourage progress by admonishing us to adapt even our sensations to new ideas; and it (d) allows for the universal application of the argumentative function of our language, and not only for its application within a given frame which itself can only be either described, or expressed - [1], page 34-36
@@
One of the basic assumptions of the orthodox view is that ’in our description of nature the purpose is […] to trace down, as far as it is possible, relations between the manifold aspects of our experience’. […] If we add to this the idea that ’only with the help of classical ideas is it possible to ascribe an unambiguous meaning to the results of observation’ […] we arrive at once at the result that a microscopic theory cannot be anything but a device for the prediction of a particular kind of fact, viz. of classical states of affairs. […] It is also true that the quantum theory is the first theory of importance which to some extent satisfies the programme of Berkeley and Mach (classical states of affairs replacing the ’perceptions’ of the former and the ’elements’ of the latter). But it must not be forgotten that there is a whole tradition which is connected with the philosophical position of realism and which went along completely different lines. In this tradition the facts of experience, whether or not they are now describable in terms of a universal theory (such as classical mechanics), are not regarded as unalterable building stones of knowledge; they are regarded as capable of analysis, of improvement, and it is even assumed that such an analysis and improvement is absolutely necessary. [examples: Galileo and Copernico, atomic theory] […] This tradition proceeds from the very reasonable assumption that our ideas as well as our experiences may be erroneous and that the latter gives at most an approximate account of what is going on in reality. Bohm’s own point of view is closely connected with this tradition. Having shown that all the attempts to prove the uniqueness of the Copenhagen Interpretation are invalid, he suggests ’to take the field and particle concepts of classical physics as starting points and to modify and enrich them in such a way that they are able to deal with the new combination of wave and particle properties that is implied in the quantum theory’ (98, my italics). Such modified concepts, or even a completely new conceptual apparatus which does not any longer make use of classical ideas, will of course at first be ’extraphysical’ (99) in the sense that it will not be accessible to test with the help of methods available before it was conceived. However, ’the history of scientific research is full of examples in which it was very fruitful indeed to assume that certain objects and elements might be real, long before any procedures were known that could permit them to be observed directly. […] In this way positivism ’constitutes a dogmatic restriction of the possible forms of future experience’ […]’ - [1], page 224-225
@@
Here we see PKF’s preference for the realist view associated with Bohmian mechanics, stressing how the latter had to incorporate the lessons of QM and modify the concepts of classical physics to account for them, even while using the same concepts as the basis for a realist position. The new concepts are adopted because a realist position is maintained and this position forces us to question the immediate results of observations, and the concepts initially used to interpret them. An instrumentalist (or ”positivist”) attitude, PKF argues, makes this critical analysis of observational concepts more difficult, and ends up with “a dogmatic restriction of the possible forms of future experience”.
@@
So, realism is always to be preferred to instrumentalism, as a more fruitful (because less dogmatic) attitude, thus for heuristic reasons, rather than a priori philosophical arguments. It leads to questioning more our assumptions, it includes the inputs from experience among the aspects of our scientific understanding of the world that require critical scrutiny, and it is supported by the realization that also our observations are loaded with theoretical elements. Let us analyse this strong conclusion further.
@@
First of all, the basic point here is the attention by PKF to what scientists actually do in practice and to which philosophical standpoint actually captures or helps in their practice. For PKF, choosing between instrumentalism and realism means choosing a philosophical interpretation (of the relationship between theory and reality) that takes seriously the actual practice of science, i.e. what scientist do when they apply the theory in concrete experimental situations. This is a crucial aspect of most of PKF’s production, in fact (in the context of PKF work on QM it has been noticed and analysed, for example, in [26]. In this way, we argue, instrumentalism and realism, and the choice between them, become (parts of) methodological rather than metaphysical positions. This happens naturally, for several reasons. Actual scientific practice is, for PKF, the main basis as well as testing ground for any methodological position, rather than any abstract principle of rationality. A good methodology is one that takes on board all the elements that help, in a given context, advancing science. This is at the descriptive level but alo at the normative level. Thus, if scientist do make progress by adopting a given interpretation, this is a main argument in favor of that interpretation. Also, while the specific interpretative framework one adopts may not directly influence the laboratory practices one uses (in this sense it may sound misplaced to see the adoption of the framework as part of a methodology), sometimes it does, focusing attention on one or another phenomenon or specific techniques. A given interpretation may also make it easier to consider new theoretical hypotheses, to modify the same theoretical framework or appreciate the value of a different one. In this sense, flexibility at the interpretative level (even leaving untouched laboratory practices) can facilitate the application of key methodological principles and associated practical steps, like the principle of proliferation, so dear to PKF. It is recognised (and PKF did at several points) that empirical assessment is only part of theory assessment (although the most solid and decisive one) and laboratory practices only part of theory development, while conceptual shifts and interpretative frameworks can play substantial role in both, so considerations about them are inevitably about methodology as well. The key aspect to be appreciated in PKF’s analysis is thus the very fact that realism is understood here as a methodological position, and supported as such, rather than a metaphysical position. It takes seriously, at the methodological level, but does not commit to, at the metaphysical level, the representational aspects of physical theories and the role they play in scientific practice.
@@
Second, we suggest that, by turning realism into a methodological position, PKF is anticipating R. Giere’s constructive realism (emphasizing the nature of scientific theories as human constructions as well as recognizing the absence of sharp distinction between theoretical and observational elements in them) [27], but even more his proposed strategy of promoting any metaphysical presupposition into a methodological maxim, and then judging the metaphysical presupposition on the basis of its fruitfulness once translated into a methodological prescription [28]. This strategy, which appears to us entirely analogous to what PKF is proposing for realism in general, is a key element in Giere’s formulation of perspectival realism [28] and it is also adopted by Giere in advocating for his more general programme of a ”naturalised philosophy of science”, where it is applied to the very notion of naturalism [29]. The resonances between PKF’s philosophy and Giere’s have been also pointed out by Brown in [30], as well as by Giere himself [31], although mostly referring to PKF’s later work. We point out that such resonances can be already found in his earlier production. This ‘perspectival’aspect of realism [^8] On the perspectival aspects and PKF’s general stance on Realism see also Chang in [5], apart from being a fertile and growing research direction in general philosophy of science [32], it has been applied by several authors to the interpretation of QM, mostly in relation to neo-Copenhagen perspectives (see e.g. [33, 34]).

### v3

@@ 3 Realism vs instrumentalism, in the interpretation of QM
PKF’s target is the orthodoxy of the time, and its instrumentalist and positivist core. We note here, though, that we could direct a similar criticism, and with the same philosophical arguments, toward any realist interpretation of QM in which the physical reality we decide to preserve, despite the challenges of QM, corresponds to an ontology built on the same premises of our observational language or our common sense, that is, upon reflection, on classical Newtonian mechanics. An example would be the same deBroglie-Bohm theory that was the only contender of the orthodox Copenhagen view at the time of PKF’s writing. The criticism would then extend to more modern perspectives, Bohmian or not, aiming at re-establishing a primitive ontology of particles moving in space and in presence of an absolute time, an ontology which is essentially Newtonian [32], at the heart of fundamental physics. Relatedly, it is interesting to look at modern ”hidden variables” completions of QM, formalized in the mathematical framework of ”ontological models” [33], where quantum states are epistemic states (states of knowledge) over a hidden (from QM) set of ontic states of the physical system under consideration. These models (which at present can cover only ”fragments” of QM) are indeed forced by the lessons of QM (including both theoretical developments of the last decades and its observational successes) to look for a precise characterization of the ontic structure of the world (what is real) in terms much farther away from classical realism and a Newtonian picture than the early advocates of a realist alternative to the Copenhagen instrumentalism (and the Bohmians) would have liked (see, e.g. [34]).
@@
It is instructive, then, to read further PKF’s critical position about this dogmatic turn of empiricism, leading to the instrumentalist Copenhagen interpretation of QM, appreciating the anti-dogmatic spirit, but also putting the realist Bohmian or hidden variable position under the same critical lens (we could try to substitute any reference to Bohr or Copenhagen in this quote with a reference to Einstein or Bohm or other ”realist” proposal), to realize that the dogmatism attacked by PKF has spread beyond the Copenhagen interpretation.
@@
The Galilean tradition, as we may call it, therefore proceeds from the very reasonable point of view that our ideas as well as our experiences (complicated experimental results included) may be erroneous, and that the latter give us at most an approximate account of what is going on in reality. Hence, within this tradition the condition to be satisfied by a future theory of the microcosm is not that it be simply compatible with duality and the other laws used in the above argument, but that it be compatible with duality to a certain degree of approximation which will have to depend on the precision of the experiments used for establishing the ’fact’ of duality. - [1], page 316-325
@@
In fact, conceptual revision is possible even about experimental facts which are taken to be exact and perfectly valid, when the concepts to be revised are strongly entrenched in everyday thinking, and when the new concepts to be introduced (which are in fact invented to establish a new theoretical paradigm) are incompatible with the existing ones. The historic example is again taken from Galileo’s story:
@@
[…] How else could it have been possible […] to replace the Aristotelian physics and the Aristotelian cosmology by the new physics of Galileo and Newton? The only conceptual apparatus then available was the Aristotelian theory of change with its opposition of actual and potential properties, form and matter, the four causes and the like. This conceptual apparatus was much more general and universal than the physical theories of today for it contained a general theory of change, spatiotemporal and otherwise. It also seems to be closer to everyday thinking and was therefore more firmly entrenched than any succeeding physical theory, classical physics included. Within this tremendously involved conceptual scheme Galileo’s (or rather Descartes’) law of inertia does not make sense. Should, then, Galileo have tried to get on with the Aristotelian concepts as well as possible because these concepts were the only ones in actual use and as ’there is no use discussing what could be done if we were other [i.e. more ingenious] beings that we are?’ By no means! What was needed was not improvement, or delimitation of the Aristotelian concepts to ’make room for new physical laws’; what was needed was an entirely new theory. Now at the time of Galileo human beings were apparently able to do this extraordinary thing and become ’beings different from what they were before’ (and one should again realize that the conceptual change that was implied was much more radical than the conceptual change necessitated by the appearance of the quantum of action). Are there (apart from pessimism with respect to the abilities of contemporary physicists) any reasons to assume that what was possible in the sixteenth and seventeenth centuries will be impossible in the twentieth century? - [1], page 316-325
@@
This result is exactly as it should be. Any restrictive demand with respect to the form and the properties of future theories can be justified only if an assertion is made to the effect that certain parts of the knowledge we possess are absolute and irrevocable. Dogmatism, however, should be alien to the spirit of scientific research, and this quite irrespective of whether it is now grounded upon ’experience’ or upon a different and more ‘aprioristic’ kind of argument. - [1], page 316-325
@@
Having established that the distinction between theories and (empirical) facts is not an absolute one and that the observational language is a theoretical one, PKF argues that the choice between realism and instrumentalism amounts to a strategic (and heuristic) choice in our pursue of a scientific understanding of nature. It has to do with which concepts (whether those grounding the (more) theoretical language or those grounding the (more) observational one) we decide to rely on when trying to improve our scientific understanding of the world (see also [35]). As with any strategic decision, we should always remain flexible and constantly adjust our strategies as we move forward and question the assumptions at the basis of them. We should, in other words, be ‘rationally tightrope-walking’.
@@
In other words, preferring realism over instrumentalism, or vice versa, is not something that can be decided on philosophical grounds (especially philosophical arguments regarding theory comparison, which are hindered by incommensurability), nor on empirical grounds, but by its results in facilitating progress, on a case-by-case and tentative basis, even though the concrete implementation of a new general perspective can be difficult. And the fact that one cannot decide a priori in favor of one option over the other does not make the choice an arbitrary one either, because the consequences for scientific progress of taking one option or the other can be judged. One could say that heuristic value is a more important criterion than epistemological or metaphysical basis (on this point, see [36]). In PKF’s words:
@@
The [second] objection is that different ideals of knowledge cannot be realized equally easily. Taking the second objection first, we are prepared to admit that there may be psychological difficulties in inventing theories of a certain type, especially if metaphysical views are held which seem to recommend radically different theories. However, a stronger variant of the second objection is frequently used. According to this stronger variant, all our theoretical knowledge is (uniquely) determined by the facts and cannot be chosen at will. Against this objection we repeat that what is determined by the ’facts’ is the acceptance (or rejection) of sentences which are already interpreted and which have been interpreted independently of the phenomenological character of what is observed. […] such a situation arises whenever a fairly general point of view was held long enough to influence our expectations, our language and thereby our perceptions, and when during that period no alternative picture was ever seriously considered. We may prolong such a situation either by explaining away adverse facts with the help of ad hoc hypotheses which are framed in terms of the points of view to be conserved; or by reducing more successful alternatives to ’instruments of prediction’ which, being devoid of descriptive meaning, cannot clash (in the phenomenological sense) with any experience; or by devising a criterion of significance according to which such alternatives are meaningless. […] But we can also choose the opposite procedure, i.e. we can take refutations seriously and regard alternative theories, in spite of their unusual character, as descriptive of really existing things, properties, relations etc. In short, although the truth of a theory may not depend upon us, its form (and the form of our theoretical knowledge in general) can always be arranged so as to satisfy certain demands. - [1], page 34-36
@@
Once more, the only real enemy is dogmatism, in all its forms and critical (self-)assessment is mandatory and should be a constant element of our investigations, whether we initially decide to proceed within a realist perspective or from an instrumentalist standpoint (again, a similar understanding is expressed in [18]. One has to exercise tightrope-walking rationality, avoiding both a-critical realism and hypocritical instrumentalism, and keep going.
@@
Having accepted this fundamentally anti-dogmatic stance, does PKF have arguments, however tentative and purely strategic in nature, for one or the other option? He does, and can be seen already from the last part of the previous quote, in which it is suggested to take a realist stance on alternative theories, even when blatantly in contrast with established ’facts’, to check their heuristic value. It can be argued, he maintains, that realism tends to be less dogmatic than instrumentalism, since it does not take even empirical facts and observations for granted (for a discussion of this point, see [37]). Indeed, it treats all elements of scientific theories as subject to critical analysis, and leaves more room for metaphysical speculation. In this sense, it is always to be preferred [^5] It is interesting to notice that, starting from the same anti-dogmatic stance, and in fact based on a similar emphasis on the existence of alternative explanations for the same phenomena, one can also be led to prefer an instrumentalist rather than realist attitude. This is the case, for example, of K. Stanford’s version of instrumentalism [38].. This is a strong conclusion.
@@
Here are PKF’s words [^6] Incidentally, the following quote is a justification of creativity in theory construction unhinged by empirical constraints and of non-empirical theory assessment, and a reminder that empirical evidence remains the ultimate judge of of our theories.:
@@
As opposed to positivism, a realistic position does not admit any dogmatic and incorrigible statement into the field of knowledge. Hence, also, our knowledge of what is observed is not regarded as unalterable and this in spite of the fact that it may have a counterpart in the phenomena themselves. This means that at times interpretations will have to be considered which do not ’fit’ the phenomena and which clash with what is immediately given. Interpretations of this kind could not possibly emerge from close attention to the ’facts’. It follows that we need a non-observational source for interpretations. Such a source is provided by (metaphysical) speculation which is thus shown to play an important role within realism. However, the results of such speculation must be made testable, and having been transformed in this way they must be interpreted as descriptive of general features of the world (otherwise we are thrown back upon the old account of what is observed). This procedure (a) allows us to draw a clear boundary line between objective states of affair and the states of the observer, though it admits that we may be mistaken with regard to the exact position of the boundary line; it (b) is empirical in the sense that no dogmatic statements are allowed to become elements of knowledge; it (c) is liable to encourage progress by admonishing us to adapt even our sensations to new ideas; and it (d) allows for the universal application of the argumentative function of our language, and not only for its application within a given frame which itself can only be either described, or expressed - [1], page 34-36
@@
[…] It is also true that the quantum theory is the first theory of importance which to some extent satisfies the programme of Berkeley and Mach (classical states of affairs replacing the ’perceptions’ of the former and the ’elements’ of the latter). But it must not be forgotten that there is a whole tradition which is connected with the philosophical position of realism and which went along completely different lines. In this tradition the facts of experience, whether or not they are now describable in terms of a universal theory (such as classical mechanics), are not regarded as unalterable building stones of knowledge; they are regarded as capable of analysis, of improvement, and it is even assumed that such an analysis and improvement is absolutely necessary. [examples: Galileo and Copernico, atomic theory] […] This tradition proceeds from the very reasonable assumption that our ideas as well as our experiences may be erroneous and that the latter gives at most an approximate account of what is going on in reality. Bohm’s own point of view is closely connected with this tradition. - [1], page 224-225
@@
Here we see PKF’s preference for the realist view associated with Bohmian mechanics (quoting directly from Bohm), stressing how the latter had to incorporate the lessons of QM and modify the concepts of classical physics to account for them, even while using the same concepts as the basis for a realist position. The new concepts are adopted because a realist position is maintained and this position forces us to question the immediate results of observations, and the concepts initially used to interpret them. An instrumentalist (or ”positivist”) attitude, PKF argues, makes this critical analysis of observational concepts more difficult, and ends up with “a dogmatic restriction of the possible forms of future experience” (in Bohm’s own words, as quoted in [1], page 224-225).
@@
So, realism is always to be preferred to instrumentalism, as a more fruitful (because less dogmatic) attitude for heuristic reasons, rather than a priori philosophical arguments. It leads to questioning more our assumptions, it includes the inputs from experience among aspects of our scientific understanding of the world that require critical scrutiny, and it is supported by the realization that also our observations are loaded with theoretical elements. Let us analyse this strong conclusion further.
@@
First, for PKF, choosing between instrumentalism and realism means choosing a philosophical interpretation (of the relationship between theory and reality) that takes seriously the actual practice of science, i.e. what scientist do when they apply the theory in concrete experimental situations. This is a crucial aspect of most of PKF’s production, in fact (in the context of PKF work on QM it has been noticed and analysed, for example, in [39]). In this way, we argue, instrumentalism and realism, and the choice between them, become (parts of) methodological rather than metaphysical positions. This happens naturally, for several reasons. Actual scientific practice is, for PKF, the main basis as well as testing ground for any methodological position, rather than any abstract principle of rationality. A good methodology is one that takes on board all the elements that help, in a given context, advancing science. This is at the descriptive level but alo at the normative level. Thus, if scientist do make progress by adopting a given interpretation, this is a main argument in favor of that interpretation. Also, while the specific interpretative framework one adopts may not directly influence the laboratory practices one uses (in this sense it may sound misplaced to see the adoption of the framework as part of a methodology), sometimes it does, focusing attention on one or another phenomenon or specific techniques. A given interpretation may also make it easier to consider new theoretical hypotheses, to modify the same theoretical framework or appreciate the value of a different one. In this sense, flexibility at the interpretative level (even leaving laboratory practices untouched) can facilitate the application of key methodological principles and associated practical steps, like the principle of proliferation, so dear to PKF. PKF recognised at several points that empirical assessment is only part of theory assessment (although the most solid and decisive one) and laboratory practices only part of theory development, while conceptual shifts and interpretative frameworks can play substantial role in both, so considerations about them are inevitably about methodology as well (see [40] for a related discussion). The key aspect to be appreciated in PKF’s analysis is that realism is understood as a methodological position, and is supported as such, rather than a metaphysical position. It takes seriously, at the methodological level, but does not commit to, at the metaphysical level, the representational aspects of physical theories and the role they play in scientific practice.
@@
Second, we suggest that, by turning realism into a methodological position, PKF is anticipating R. Giere’s constructive realism (emphasizing the nature of scientific theories as human constructions as well as recognizing the absence of sharp distinction between theoretical and observational elements in them) [41], but even more his proposed strategy of promoting any metaphysical presupposition into a methodological maxim, and then judging the metaphysical presupposition on the basis of its fruitfulness once translated into a methodological prescription [42]. This strategy, which appears to be entirely analogous to what PKF is proposing for realism in general, is a key element in Giere’s formulation of perspectival realism [42] and it is also adopted by Giere in advocating for his more general programme of a ”naturalised philosophy of science”, where it is applied to the very notion of naturalism [43]. The resonances between PKF’s philosophy and Giere’s have been also pointed out by Brown in [44], as well as by Giere himself [45], although mostly referring to PKF’s later work. We point out that such resonances can be found in his earlier production. This ‘perspectival’ aspect of realism [^7] On the perspectival aspects and PKF’s general stance on Realism see also Chang in [5], apart from being a fertile and growing research direction in general philosophy of science [46], it has been applied by several authors to the interpretation of QM, mostly in relation to neo-Copenhagen perspectives (see e.g. [47, 48]).